# SN 2024abvb: A Type Icn Supernova in the Outskirts of its Host Galaxy

### v1

@@ IV Discussion and Conclusions
In this paper, we present photometry and spectroscopy of SN 2024abvb, which we classify as an SN Icn based on the identification of a set of relatively narrow (FWHM –2000 km s-1), singly ionized carbon emission lines. Such a carbon-rich CSM expanding at a fairly high velocity implies a hydrogen- and helium-depleted, carbon-dominant, Wolf-Rayet-star-like wind (WC type). Among the SN Icn subclass, the light-curve morphology exhibits significant diversity with an -band peak magnitude ranging from 17 mag (SN 2019jc, 2022ApJ...938...73P) to 20 mag (SN 2021csp, 2022ApJ...927..180P; 2022ApJ...938...73P). SN 2024abvb shows a peak brightness similar to that of SN 2021csp but declines faster.
We fit the UV-to-optical light curves of SN 2024abvb, adopting a hybrid analytic model that accounts for both the ejecta-CSM interaction and the radioactive decay of 56Ni. The best-fit result suggests a rather small nickle mass of M⊙ in the ejecta, and the near-peak emission of SN 2024abvb is dominated by the ejecta-CSM interaction. This low Ni mass is consistent with the analysis of other SNe Icn.
@@
The nature of SNe Icn still remains controversial. Similar to SNe Ibn, they may originate from the explosion of WR stars, since both exhibit signatures of CSM that expand at rather high velocities of and have stripped their outer H/He envelopes. The mechanism driving such intense ejection of the outer H/He envelopes immediately prior to the terminal explosion remains unexplained. The absence of apparent P Cygni profiles in our early-time spectral sequence of SN 2024abvb hinders the measurement of the wind velocity. The modeled () and () are smaller than those of other SNe Icn or stripped-envelope SNe as shown in Figure 9, indicating that the progenitor star of SN 2024abvb experienced significant mass loss prior to the explosion.
@@
SN 2024abvb exploded in the outskirts of its host galaxy, which may correspond to a relatively long evolving history for its progenitor system. One plausible explanation for the small ejecta mass and nickel mass associated with the bright, fast-declining light curve of SN 2024abvb is that the exploding progenitor star originated in a binary system consisting of a helium star and a compact companion (2015MNRAS.451.2123T). The H and He envelopes of the former are progressively stripped by the latter through Roche-lobe overflow (2007ASPC..372..397M). Such enhanced mass loss toward the orbital plane of the binary system would take place when the Roche lobe is filled, and is suggested to be capable of efficiently stripping the H envelope (2014ARA&A..52..487S).
@@
The chemical composition and wind velocity of the CSM encode information about the mass-loss history of massive stars nearing the end of their lives. The fusion reactions transform massive stars into a structure of concentric shells. One of the key deliverables of stripped-envelope SNe is testing the long-hypothesized internal layering that develops at the end of the evolution of massive stars: progressively heavier elements are synthesized toward the iron core at the center. Spectrophotometric observations from the earliest possible phases would reveal the composition and physical conditions of the surface layers of the exploding star, which are immediately expelled before the explosion. Observations of Type Ibn/Icn SNe, together with the very recently reported Type Ien SN 2021yfj (2024TNSAN.239....1G; 2024TNSAN.240....1G; 2024arXiv240902054S), whose immediate CSM exhibits chemical signatures of silicon, sulfur, and argon, provide key clues for understanding the nature and process of such a stripping sequence emerging from the death of massive stars.

### v2

@@ IV Discussion and Conclusions
In this paper, we present photometry and spectroscopy of SN 2024abvb, which we classify as an SN Icn based on the identification of a set of relatively narrow (FWHM –2000 km s-1), singly ionized carbon emission lines. Such a carbon-rich CSM expanding at a fairly high velocity implies a hydrogen- and helium-depleted, carbon-dominant, Wolf-Rayet-star-like wind (WC type). Among the SN Icn subclass, the light-curve morphology exhibits significant diversity with an -band peak magnitude ranging from 17 mag (SN 2019jc, Pellegrino et al., 2022) to 20 mag (SN 2021csp, Perley et al., 2022; Pellegrino et al., 2022). SN 2024abvb shows a peak brightness similar to that of SN 2021csp but declines faster.
We fit the UV-to-optical light curves of SN 2024abvb, adopting a hybrid analytic model that accounts for both the ejecta-CSM interaction and the radioactive decay of 56Ni. The best-fit result suggests a rather small nickel mass of M⊙ in the ejecta, and the near-peak emission of SN 2024abvb is dominated by the ejecta-CSM interaction. The derived explosion time, kinetic energy, 56Ni mass, and CSM mass are consistent with the results in Aster et al. (2026), although they obtained a relatively larger ejecta mass (). This difference may arise from differing assumptions about the CSM: we adopted a steady-wind CSM model, whereas Aster et al. (2026) used a shell-like CSM model. Besides, Shi et al. (2026) fitted the multi-band light curve of SN 2024abvb to constrain the 56Ni mass of , indicating the potential 56Ni-driven picture. The relatively large kinetic energy () in Shi et al. (2026) produces an expansion velocity of about (), possibly resulting in the fast-evolving light curve as SN 2024abvb does. However, the lack of a late-phase light curve for SN 2024abvb hinders the analysis of the synthesized 56Ni, as the energy contribution from 56Co 56Fe increases and the contribution from CSM interaction decreases.
@@
The nature of SNe Icn still remains controversial. Similar to SNe Ibn, they may originate from the explosion of WR stars, since both exhibit signatures of CSM that expand at rather high velocities of and have stripped their outer H/He envelopes. The mechanism driving such intense ejection of the outer H/He envelopes immediately prior to the terminal explosion remains unexplained. The absence of apparent P Cygni profiles in our early-time spectral sequence of SN 2024abvb hinders the measurement of the wind velocity. The modeled () and () are smaller than those of stripped-envelope SNe as shown in Figure 9, indicating that the progenitor star of SN 2024abvb experienced significant mass loss prior to the explosion.
In the – relationship shown in Figure 9, it seems that SN 2024abvb might be consistent with the ultrastripped-envelope SN (USSN) group, indicating that they may share a similar physical origin. The key difference lies in the presence of dense confined CSM around SN 2024abvb, whereas USSNe exhibit a relatively clean circumstellar environment. This distinction accounts for the higher luminosity of SN 2024abvb compared to USSNe (e.g., -band peak magnitude for SN 2021agco Yan et al. 2023). If the underlying explosions are indeed the same, SN 2024abvb should exhibit nebular-phase spectral features comparable to those of USSNe. With the increasing number of time-domain surveys, the physical origins linking SNe Icn to USSNe will be further constrained.
@@
SN 2024abvb exploded in the outskirts of its host galaxy, which may correspond to a relatively long evolving history for its progenitor system. One plausible explanation for the small ejecta mass and nickel mass associated with the bright, fast-declining light curve of SN 2024abvb is that the exploding progenitor star originated in a binary system consisting of a helium star and a compact companion (Tauris et al., 2015). The H and He envelopes of the former are progressively stripped by the latter through Roche-lobe overflow (Mohamed and Podsiadlowski, 2007). Such enhanced mass loss toward the orbital plane of the binary system would take place when the Roche lobe is filled, and is suggested to be capable of efficiently stripping the H envelope (Smith, 2014). The INTEL Collaboration et al. (2026) reported the time-variant polarimetric signals and the presence of Balmer absorptions and He lines in their high-resolution spectra. These observations suggest that SN 2024abvb arises from a binary system undergoing multiple mass-loss episodes with different orientations during the common-envelope phase. While this scenario adds further details, it fundamentally aligns with the physical picture described above.
@@
The chemical composition and wind velocity of the CSM encode information about the mass-loss history of massive stars nearing the end of their lives. The fusion reactions transform massive stars into a structure of concentric shells. One of the key deliverables of stripped-envelope SNe is testing the long-hypothesized internal layering that develops at the end of the evolution of massive stars: progressively heavier elements are synthesized toward the iron core at the center. Spectrophotometric observations from the earliest possible phases would reveal the composition and physical conditions of the surface layers of the exploding star, which are immediately expelled before the explosion. Observations of Type Ibn/Icn SNe, together with the very recently reported Type Ien SN 2021yfj (Gal-Yam et al., 2024b, a; Schulze et al., 2024), whose immediate CSM exhibits chemical signatures of silicon, sulfur, and argon, provide key clues for understanding the nature and process of such a stripping sequence emerging from the death of massive stars.